OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). sin^2 x + sin x = 0

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \sin^2x+\sin x = 0}\) \(\color{#000000 }{ \displaystyle \sin x(\sin x +1)= 0}\) will divide both sides by \(\color{#000000 }{ \displaystyle \sin x }\), and we get. \(\color{#000000 }{ \displaystyle \sin x+1 = 0}\) and note that when we divide by \(\color{#000000 }{ \displaystyle \sin x }\), then \(\color{#000000 }{ \displaystyle \sin x = 0}\) is also a solution.

OpenStudy (solomonzelman):

So you are going to have to equations to solve, \(\color{#000000 }{ \displaystyle \sin x = 0}\) and \(\color{#000000 }{ \displaystyle \sin x +1= 0}\)

OpenStudy (anonymous):

Okay that makes sense. How would I go about solving these equations?

OpenStudy (solomonzelman):

Well, you can refer to the unit circle, and tell me for what values of x, will sin(x) be equal to zero?

OpenStudy (anonymous):

pi?

OpenStudy (anonymous):

Then for the next one it would be sin x = -1 so the answer would be 3pi/2?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Yes, for sin(x)=0, x=π, and another solution is x=0. (0 is in the interval) And for sin(x)=-1 you also got the correct solution, x=3π/2

OpenStudy (anonymous):

Okay, thank you very much!